-4(3+x)+5=4(x+3).....can i get a detailed explenation of how to do this problem, i dont know if i should add 4x to each side or not after using the distributive property.
Answer by checkley77(12844)

-4(3+x)+5=4(x+3)
-12-4x+5=4x+12
-4x-4x=12+12-5
-8x=19
x=19/-8
x=-19/8 ans.
Proof:
-4(3-19/5)+5=4(-19/8+3)
-12+76/8+5=-76/8+12
76/8+76/8=12+12-5
152/8=19
19=19